the telephone extensions at a company use 4 digits

a. how many extensions are possible if there no restrictions?
b. how mant extensions are possible if the first digit cannot be 0 or 9?
c. how many extensions are possible if the first digit can only be 1?

Found 2 solutions by stanbon, checkley77:
Answer by stanbon(75887)

the telephone extensions at a company use 4 digits
a. how many extensions are possible if there no restrictions?
Ans: 10^4 = 10,000
-----------------------
b. how mant extensions are possible if the first digit cannot be 0 or 9?
Ans: 8*10^3 = 8000
----------------------
c. how many extensions are possible if the first digit can only be 1?
Ans: 1*10^3 = 1000
========================
Cheers,
Stan H.

Answer by checkley77(12844)

10^4=10,000 POSSIBLE 4 DIGIT EXTENSIONS.
8*10^3=8,000 POSSIBLE EXTENSIONS WITHOUT A 0 OR 9 AS THE FIRST DIGIT.
1*10^3=1,000 POSSIBLE EXTENSIONS BEGINNING WITH A 1.